OpenStudy (anonymous):

math question of the day: Prove that \[\forall n \in \mathbb{N} (n \ge 1 \rightarrow n^{n} \ge n)\]

OpenStudy (anonymous):

oopes thats \[n^{n} \ge n!\]

OpenStudy (hoblos):

i think it can be solved by induction....

OpenStudy (hoblos):

for n=1 1 >=1 true

OpenStudy (hoblos):

if we suppose it is true for n n^n >= n!

OpenStudy (nenadmatematika):

then we proove that is correct for n+1 so (n+1)^(n+1)>=(n+1)! (n+1)^n*(n+1)>=(n+1)*n! after cancelling n+1 we get (n+1)^n>=n! so our assumption was that n^n>=n! so we proved it...is this OK hoblos ?

OpenStudy (hoblos):

for n+1 (n+1)^(n+1) = (n+1)(n+1)^n but we know that (n+1)>n (n+1)^n >n^n since n>=1 then (n+1)^(n+1) = (n+1)(n+1)^n > (n+1)n^n > (n+1)*n! > (n+1)!

OpenStudy (hoblos):

@nenadmatematika i think your way is right too